# Non-coherent evolution of closed weakly interacting system leads to equidistribution of probabilities of microstates

### v4

@@ Abstract
We introduce a concept of non-coherent evolution of macroscopic quantum systems. We show that for weakly interacting systems such evolution is a Markovian stochastic process. The transition rates between system states, which characterize the process, are determined by Fermi’s golden rule. Such evolution is time-irreversible and leads to the equidistribution of probabilities across every state of the system. Furthermore, we investigate the time dependence of the mean numbers of particlesin single-particle states and find that, under the given assumptions, it is governed by the Boltzmann collision integral. In this theory, non-coherence is the mechanism that transforms time-reversible unitary evolution into time-irreversible stochastic evolution. Thus, we present a possible solution to the famous arrow-of-time problem.
@@ I Introduction
In the textbook [14], the second law of thermodynamics is derived from an equation known as the Fermi master equation. However, it is noted that the equation itself lacks a proper derivation. A development of this idea is the Lindblad master equation [15, 16], which describes systems weakly coupled to their environments. These equations have been shown to exhibit irreversible behavior, even when the underlying microscopic dynamics are unitary. This approach has been applied to various models in quantum thermodynamics [17, 18]. A particularly important result from this line of research is the identification of conditions under which systems reach a steady state that mirrors the classical Boltzmann distribution [19, 20].
Modern approaches propose that irreversibility emerges from a time-asymmetric mechanism inherent in quantum mechanics [21]. Typically, decoherence is suggested as this mechanism [22]. Decoherence theory explains measurement in quantum theory as the interaction of a quantum system with its environment, which causes the loss of coherence in the system [23].
@@
In [26], the authors derive the second law of thermodynamics from a quantum-mechanical analysis of a total system partitioned into an object and its environment. This type of theory has attracted considerable attention in recent years [27, 28, 29]. Additionally, the theory of non-Markovian quantum evolution [30, 31, 32] has garnered significant interest. It focuses on cases where the system’s dynamics are influenced by interactions with the external world.
However, all of the aforementioned theories locate irreversibility within a subsystem of a larger, isolated system – the environment. The total system, being isolated, is still governed by reversible equations, which represents a significant conceptual gap [35]. In the classical Gibbs treatment of subsystems, they are considered part of a larger system in equilibrium, described by the microcanonical ensemble. Reversible equations, however, cannot reach this stable state [1]. From the perspective of the second law of thermodynamics, satisfying the condition of entropy growth for the entire system requires postulating a low-entropy initial state. But this is the solution, which the authors of the cited works dismiss as ad hoc and unsatisfactory.
To address this, in the present paper we use the concept of decoherence but approach the notion of non-coherence from a different perspective than decoherence theory. Rather than focusing on the loss of coherence in a quantum system due to interaction with its environment, our investigation centers on the environment itself. We treat it as inherently non-coherent. We observe that it is the coherent states of matter that are exotic, while non-coherent states occur naturally. Coherent states – such as superconductive and superfluid phases – arise only in specific substances under low temperatures. Coherent light is generated by lasers. In contrast, natural light is typically non-coherent. From this, we infer that non-coherence represents the inherent, natural state of matter.
@@
The most significant difference between the approach presented here and decoherence theory is that the latter, in the spirit of the original Copenhagen interpretation, divides changes occurring in a quantum system into two categories: normal evolution, described by the Schrödinger equation, and measurement. The latter is then associated with the process of decoherence. Of these two parts, it is measurement that is considered irreversible [22]. According to this theory, decoherence arises due to the interaction of the quantum system with its environment. However, it is not universally accepted that decoherence theory resolves the measurement problem [35]. We treat decoherence differently, drawing on an analogy with wave optics. In optics, non-coherence arises because no wave with a continuous spectrum is entirely monochromatic. As a result, light emitted from different sources cannot form stable interference patterns, and waves emitted from the same source will desynchronize if split and travel along paths whose length difference exceeds the coherence length [36]. In our view, it is not the process of coherence loss that is irreversible, but rather the evolution of a system large enough that different parts of it are no longer mutually coherent.
We aim to comprehend how a system described by quantum theory but lacking coherence evolves. The mathematical procedure used to derive equations for non-coherent light involves averaging the arguments of the complex amplitudes of waves. A parallel between quantum entanglement and the coherence of light was recently observed in Ref. [37]. Our objective is to extend the procedure of averaging the arguments of complex amplitudes to general quantum systems and derive the equation governing non-coherent evolution.
@@ II Non-coherent transformations
As stated in the introduction, the concept of non-coherent transformation was introduced in Ref. [33] to describe the transmission of quantum particles through an interface. It was proposed to explain the emergence of irreversibility in such cases. As a physical example of the formalism presented below, one can imagine a beam of quantum particles incident on an interface and being scattered in different directions. However, except for one important case, we will not refer to this specific example and will proceed in an abstract manner.
@@
[…]
@@
Figure 1: Two beams of light are incident on the interface and are transformed either coherently or non-coherently, depending on the conditions. The interface can be, for example, between air and optical glass (depicted as light blue). The gray rectangles represent mirrors. In sub-pictures (a) and (b), two beams are emitted by the same source and separated by the interface. The initial state of the system is defined by the amplitudes after separation. The column of initial amplitudes is denoted as $\mathcal{A}_{i}$, and the column of initial intensities as $\mathcal{P}_{i}$.
If the difference between the optical paths of the two split beams is smaller than the coherence length of the light, $\Delta l<l_{\text{coh}}$, a coherent transformation occurs according to formula (3), as depicted in sub-picture (a). This transformation is reversible, and with appropriate phase differences between the beams, they can merge back into a single beam, meaning the splitting transformation has been reverted. If the difference between the optical paths exceeds the coherence length of the light, $\Delta l>l_{\text{coh}}$, a non-coherent transformation occurs according to formula (6). The resulting intensities no longer depend on the phase difference and are strictly greater than zero, which means the split beams cannot merge back into a single beam. In sub-picture (c), two beams from different sources meet at the interface. They cannot form a stable interference pattern, which means they transform irreversibly, in accordance with formula (6).
@@
In Ref. [34], an optical experiment is conducted to demonstrate the irreversibility of non-coherent transmission of light through the interface. First, a beam of light is split into two parts by the interface between air and optical glass. Then, after reflecting off mirrors, the two parts meet again at the interface (Fig. (1)a, b). If the difference in the optical path lengths exceeds the coherence length, $\Delta l>l_{\text{coh}}$, the resulting intensities of the beams can be calculated using a product of bistochastic matrices, without considering the phase difference between the beams. In this case, formula (6) should be used. However, if the difference in the optical path is smaller than the coherence length, $\Delta l<l_{\text{coh}}$, the actual phase difference between the two parts of the light must be taken into account, so we calculate using formula (3). Importantly, in the latter case, if the phase difference between the two beams differs by $2\pi n$, the beams reunite into one at the interface—demonstrating the reversibility of the process.
@@
Let us abstract away the details of the described experiment and formulate the mathematical essence. The formulas we provide in the following few paragraphs are well-known in the contexts of classical [36] and quantum optics [Knight], so we will not provide extensive detail. However, we aim to rewrite them in the specific context and notation of this manuscript for further use.
@@
[…]
@@
[$&\Delta t<t_{coh},\,\text{Int}=\,1\\ &\Delta t>t_{\text{coh}},\,\text{Int}=\frac{t_{\text{coh}}}{\Delta t}\end{aligned}\right.$ (11)
@@
$I_{res}=|A_{1}|^{2}+|A_{2}|^{2}+2\,\text{Int}\cdot|A_{1}||A_{2}|\cos\phi,$ (12)
where Int is given by expression (11). We see that if $\Delta t\gg t_{coh}$, the resulting intensity is simply the sum of the intensities of the two sources. This case is referred to as the non-coherent regime.
@@
Another case where averaging over phases is applicable is when particles emitted at different points meet at the interface. If the emissions of these particles are not causally related, they cannot exhibit stable interference. The optical analog of this was studied in Refs. [38, 39], where the authors investigated interference from two different sources. While interference patterns form at any given moment, over a time interval comparable to the decoherence time, the interference pattern drifts: constructive interference may turn into destructive interference and vice versa. Therefore, when calculating the mean intensity over a time interval much greater than the coherence time, we should average over phases.
A mathematical explanation of this may be the following. If we restrict both $A_{1}$ and $A_{2}$ from formula (9) so that they are nonzero only within some finite frequency interval, the formula defines a scalar product in an infinite-dimensional Hilbert space. In $n$-dimensional space, the expected value of the scalar product of two unit vectors is proportional to $n^{-1/2}$ [40]. So as $n\to\infty$, it tends to zero. This means that if we pick two random sources, their interference defined by formula (9) is zero.
Thus, the two distinct mechanisms leading to a lack of coherence in interaction, depicted in Fig. ((1)b, (1) c), yield the same result: the correct outcome is found by averaging over phases.
We have described two mechanisms leading to the switch from coherent transformation to the non-coherent one. The first mechanism is particularly interesting because it allows for direct experimental measurements in both coherent and non-coherent regimes, and the transition between these regimes can be described theoretically by formula (LABEL:interference). However, we suppose that the second mechanism – namely, the interaction of waves whose emissions are not causally related – is the one responsible for the non-coherence of transformations of the states of macroscopic-sized matter. In later sections, we will illustrate this with concrete examples.
@@
III Non-coherent evolution equation and quantum transitions
@@
As the previous section shows, in the discrete case, the non-coherence condition, $\Delta t\gg t_{\text{coh}}$ or $\Delta E\ll\omega_{\text{width}}$, converts a reversible unitary transformation into an irreversible stochastic transformation. In the continuous case, unitary evolution is described by the exponential of a skew-Hermitian matrix [41]. In the context of quantum theory, we typically express this as $\mathcal{U}(t)=\exp(i\mathcal{H}t)$, where $\mathcal{H}$ is a Hermitian matrix, particularly representing the Hamiltonian of the system. The continuous analog of a bistochastic transformation is a symmetric continuous-time Markov chain (CTMC). The time dependence of transition probabilities is given by $\mathcal{T}(t)=\exp(\mathcal{Q}t)$, where $\mathcal{Q}$ is the transition rate matrix (Q-matrix) [42]. Its non-diagonal elements are positive numbers representing the rates of transition between states. The diagonal elements $Q_{\alpha\alpha}$ represent the total departure rate from the state $\alpha$. They are negative, such that the sums of elements in each column are zero [^1] In probability texts, the state of the system is often represented by a row, and the transformation matrix is placed on the right of it. We use the convention adopted in physical texts: to represent the state as a column and place the transformation matrix to the left of it. This convention results in a difference in definition: in probability textbooks, the sums of elements in each row of a Q-matrix are zero.. A general Q-matrix serves as the continuous analog of a stochastic transformation, while the analog of a bistochastic transformation is a Q-matrix symmetric over the main diagonal, i.e., $Q_{\alpha\beta}=Q_{\beta\alpha}$.
We are going to demonstrate that just as the unitary transformation converts into a bistochastic transformation under the condition of non-coherence, unitary evolution converts into symmetric stochastic evolution. In our consideration, we will extensively use the fact that a closed quantum system has a discrete spectrum. However, we consider a macroscopic system where a continuous spectrum with a well-defined density of states serves as a suitable approximation, which is typical for a closed macroscopic system [1]. We consider interaction in a local area of a closed system so that the emission of interacting particles would typically occur at some distance and be causally unrelated. We will specify this requirement further in Section IV after we have a good understanding of the relation of parameters necessary for the presented theory to be applicable. We also imply the condition of weak interaction, since strongly interacting systems are beyond the applicability of the methods that we use.
@@
We start similarly to a classic simple time-dependent perturbation theory [43] that is aimed to derive Fermi’s golden rule. We seek for a wave function of a many-body system in the form
@@
Equation (23) is too complicated to tackle directly. To continue our investigation, we need to use an approximate method. The original Dirac’s approach is to employ the variation of parameters method. We choose a different approach here.
@@
[…]
@@
[…]
@@
We now want to analyze the non-coherent regime. This means that the state under consideration, $|k\rangle$, belongs to a sheaf of states with a width in the frequency region $\omega_{\text{idth}}$, and we choose a working timescale, $\Delta t$, such that it is much larger than the coherence time. Equivalently, we have $\Delta E\ll\omega_{\text{idth}}$ (Fig. 2b). The quantity $\Delta E$ fits into $\omega_{\text{idth}}$ many times, each time with a different phase difference for each pair of amplitudes, so we can average over the phase. This allows the formula for a series of non-coherent transformations, given in (20), to be applicable.
@@
We begin by utilizing an approximation of a continuous spectrum. Near a given state $l$, we divide the quantum numbers into two components. The first is $l_{T}$, such that small changes in its value leave the energy unchanged, so it is tangential to the energy surface. The second is $l_{N}$, where a change results in a change in energy. We express energy as a function of the quantum numbers in the form $E(l)=E(l_{0})+l_{N},\mathrm{grad}_{l}E$. Additionally, we recall expression (27) and substitute $\Delta t=\pi\hbar/\Delta E$. We then obtain:
@@
[…]

### v5

@@ Abstract
We introduce a concept of non-coherent evolution of macroscopic quantum systems. We show that for weakly interacting systems such evolution is a Markovian stochastic process. The transition rates between system states, which characterize the process, are determined by Fermi’s golden rule. Such evolution is time-irreversible and leads to the equidistribution of probabilities across every state of the system. Furthermore, we investigate the time dependence of the mean numbers of particles in single-particle states and find that, under the given assumptions, it is governed by the Boltzmann collision integral. The proposed mechanism that transforms time-reversible unitary evolution into time-irreversible stochastic evolution is non-coherence. In the presented theory, the non-coherence is not associated with interaction with a heat bath, but rather with the finite spectral width of quantum states. This understanding of non-coherence is analogous to the one used in wave optics. Thus, we present a novel approach to the famous arrow of time problem.
@@ I Introduction
Modern approaches propose that irreversibility emerges from a time-asymmetric mechanism inherent in quantum mechanics [21]. Decoherence is typically suggested as this mechanism [22]. Decoherence theory explains measurement in quantum theory as the interaction of a quantum system with its environment, leading to the loss of coherence in the system [23]. However, it is not universally accepted that decoherence theory resolves the measurement problem [24, 25, 26]. While it does account for the vanishing of the non-diagonal elements of the density matrix, it does not explain why only one definite outcome is realized in a measurement. We will refer to this issue as the problem of outcomes.
@@
In the textbook [14], the second law of thermodynamics is derived from an equation known as the Fermi master equation. However, the equation itself is noted to lack a proper derivation. A modern justification for this idea is provided by the Lindblad master equation [15], which describes systems weakly coupled to their environments. In particular, the Fermi master equation is regarded as a Born-Markov approximation of the Lindblad equation [16]. These equations have been shown to exhibit irreversible behavior, even when the underlying microscopic dynamics are unitary. This approach has been applied to various models in quantum thermodynamics [17, 18]. A particularly important result of this line of research is the identification of conditions under which systems reach a steady state resembling the classical Boltzmann distribution [19, 20].
Analogously, in Ref. [29], the authors derive the second law of thermodynamics from a quantum-mechanical analysis of a total system partitioned into an object and its environment. This type of theory has attracted considerable attention in recent years [30, 31, 32]. Additionally, the theory of non-Markovian quantum evolution [33, 34, 35] has garnered significant interest. It focuses on cases where the system’s dynamics are influenced by interactions with the external world.
However, all of the aforementioned theories locate irreversibility within a subsystem of a larger, isolated system – the environment. The total system, being isolated, is still governed by reversible equations, which represent a significant conceptual gap [25]. In the classical Gibbs treatment of subsystems, they are considered part of a larger system in equilibrium, described by the microcanonical ensemble. Reversible equations, however, cannot reach this stable state [1]. From the perspective of the second law of thermodynamics, satisfying the condition of entropy growth for the entire system requires postulating a low-entropy initial state. But this is the solution, which the authors of the cited works dismiss as ad hoc and unsatisfactory.
To address this, in the present paper we use the concept of non-coherence but approach the notion from a different perspective than decoherence theory or quantum thermodynamics. Rather than focusing on the loss of coherence in a quantum system due to interaction with its environment, our investigation centers on the environment itself. We treat it as inherently non-coherent. We observe that it is the coherent states of matter that are exotic, while non-coherent states occur naturally. Coherent states – such as superconductive and superfluid phases – arise only in specific substances under low temperatures. Coherent light is generated by lasers. In contrast, natural light is typically non-coherent. From this, we infer that non-coherence represents the inherent, natural state of matter.
@@
The most significant difference between the approach presented here and decoherence theory is that the latter, in the spirit of the original Copenhagen interpretation, divides changes occurring in a quantum system into two categories: normal evolution, described by the Schrödinger equation, and measurement. The latter is then associated with the process of decoherence. Of these two parts, it is measurement that is considered irreversible [22]. According to this theory, decoherence arises due to the interaction of the quantum system with its environment. We treat decoherence differently, drawing on an analogy with wave optics. In optics, non-coherence arises because no wave with a continuous spectrum is entirely monochromatic. As a result, light emitted from different sources cannot form stable interference patterns, and waves emitted from the same source will desynchronize if split and travel along paths whose length difference exceeds the coherence length [38]. In our view, it is not the process of coherence loss that is irreversible, but rather the evolution of a system large enough that different parts of it are no longer mutually coherent.
The quantum thermodynamics approach [17] is the closest to the one we adopt here. However, we are not satisfied with this theory either. Decoherence as a result of interaction with an environment is a natural perspective in quantum technologies, where one seeks to create an entangled system of qubits and outside interactions destroy it. It is, however, a very unnatural perspective for the foundations of statistical physics, where one must artificially introduce a distinction between the system and the environment, since no such distinction is inherent to the problem at hand. Therefore, we aim to develop a theory without such a distinction.
Moreover, in Gibbs theory, the microcanonical ensemble, that is, the probability distribution of states for the closed system, serves as the starting point. It precedes the introduction of more specific distributions and even the definition of temperature. Furthermore, the second law of thermodynamics states that the entropy of an isolated system does not decrease. Therefore, our goal is to formulate an irreversible theory of the evolution of a closed system.
To accomplish this goal, we formulate a model of an isolated quantum system described by quantum theory but with a finite spectral width of quantum states. We aim to comprehend how a system with a finite spectral width, and hence lacking coherence, evolves. The mathematical procedure used to derive equations for non-coherent light involves averaging the arguments of the complex amplitudes of waves. Such a parallel between quantum entanglement and the coherence of light was, for example, recently observed in Ref. [40]. Our objective is to extend the procedure of averaging the arguments of complex amplitudes to general quantum systems and derive the equation governing non-coherent evolution.
@@ II Non-coherent transformations
As stated in the introduction, the concept of non-coherent transformation was introduced in Ref. [36] to describe the transmission of quantum particles through an interface. It was proposed to explain the emergence of irreversibility in such cases. As a physical example of the formalism presented below, one can imagine a beam of quantum particles incident on an interface and being scattered in different directions. However, we will mostly not refer to this specific example and will proceed in an abstract manner.
@@
We see that the integral of the sum over $k^{\prime}_{i}\neq k_{i}$ equals zero since we have $\int d\varphi_{k_{i}}\exp i(\varphi_{k^{\prime}_{i}}-\varphi_{k_{i}})=0$. The integral over the first term equals $(2\pi)^{N}\sum_{\beta}|U_{\alpha\beta}|^{2}|A_{\beta}|^{2}$, since integrated function does not depend on integration variable. We see that after averaging over arguments, the column of probabilities is
@@
Figure 1: Two beams of light are incident on the interface and are transformed either coherently or non-coherently, depending on the conditions. The interface can be, for example, between air and optical glass (shown in light blue). The gray rectangles represent mirrors. In subfigures (a) and (b), two beams are emitted by the same source and separated by the interface. The initial state of the system is defined by the amplitudes after separation. The column of initial amplitudes is denoted as $\mathcal{A}_{i}$, and the column of initial intensities as $\mathcal{P}_{i}$.
If the difference between the optical paths of the two split beams is smaller than the coherence length of the light, $\Delta l<l_{\text{coh}}$, a coherent transformation occurs according to formula (3), as depicted in subfigure (a). This transformation is reversible, and with appropriate phase differences between the beams, they can recombine into a single beam, meaning the splitting transformation has been reverted. If the difference between the optical paths exceeds the coherence length of the light, $\Delta l>l_{\text{coh}}$, a non-coherent transformation occurs according to formula (6). The resulting intensities no longer depend on the phase difference and are strictly greater than zero, which means the split beams cannot merge back into a single beam.
@@
In Ref. [37], an optical experiment is conducted to demonstrate the irreversibility of non-coherent transmission of light through the interface. First, a beam of light is split into two parts by the interface between air and optical glass. Then, after reflecting off mirrors, the two parts meet again at the interface (Fig. 1a, b). If the difference in the optical paths lengths exceeds the coherence length, $\Delta l>l_{\text{coh}}$, the resulting intensities of the beams can be calculated using a product of bistochastic matrices, without considering the phase difference between the beams. In this case, formula (6) should be used. However, if the difference in the optical path is smaller than the coherence length, $\Delta l<l_{\text{coh}}$, the actual phase difference between the two parts of the light must be taken into account, so we calculate using formula (3). Importantly, in the latter case, if the phase difference between the two beams differs by $2\pi n$, the beams reunite into one at the interface, demonstrating the reversibility of the process.
Although the experiment in [37] was carried out in the context of classical optics, the results can be interpreted in terms of photons and wave packets. First of all, the intensity of light is proportional to the number of photons multiplied by the energy flux associated with each, $c\hbar\omega$. When photons are reflected in a given direction, the expected number is given by the product of the number of incident photons and the probability of transmission, $NP$. By the law of large numbers, the actual number of photons does not differ significantly from the expected value. Thus, we can write $I=PNc\hbar\omega$. This allows us to use intensity and probability interchangeably.
Second, the coherence length corresponds to the length of the photon wave packet, and both are inversely related to its spectral dispersion. Speaking very roughly, we can subdivide the beam into packets of photons of length $l_{\text{coh}}$, and observe that if the difference in optical paths is smaller than the size of a packet, the packet will primarily interact with itself at the interface (Fig. 2a). This type of interaction results in interference and allows a description in terms of probability amplitudes $\mathcal{A}$, which are transformed by the unitary matrix $\mathcal{U}$. Otherwise, if the size of the packet is smaller than the difference in optical paths, the state interacts with another state (Fig. 2b), and such interaction does not produce interference. In this case, photons are described by probabilities $\mathcal{P}$, and the transformation of these probabilities is given by the bistochastic matrix $\mathcal{T}$.
Figure 2: As in Fig. (1), two beams of light are incident on the interface and are transformed either coherently or non-coherently, depending on the conditions. We present a qualitative scheme for interpreting the experiment in terms of photon packets. Subfigure (a) shows a beam of light divided into segments of equal length—the coherence length $l_{\text{coh}}$. Each segment roughly corresponds to a single packet of photons, and the segments are numbered sequentially from left to right.
In subfigure (b), the left end of packet number one is shown crossing the mirror. Since the difference between the optical paths of the two split parts of the beam is smaller than the coherence length, $\Delta l<l_{\text{coh}}$, the state interacts with itself, and the interaction is coherent. In particular, it is packet number one interacting with itself. In subfigure (c), the difference between the optical paths exceeds the coherence length, $\Delta l>l_{\text{coh}}$, and the state interacts with another state, resulting in non-coherent behavior. Specifically, it is state number $n$ interacting with state number $m$, with $n\neq m$.
@@
Let us abstract away the details of the described experiment and formulate the mathematical essence. The formulas we provide in the following few paragraphs are well-known in the contexts of classical [38] and quantum optics [39], so we will not provide extensive detail. However, we aim to rewrite them in the specific context and notation of this manuscript for further use.
@@
$\text{Int}=\int_{-\infty}^{\infty}d\omega A_{1}^{*}(\omega)^{*}e^{-i\omega t}A_{2}(\omega)e^{i\omega t}=\int_{-\infty}^{\infty}d\omega A_{1}(\omega)^{*}A_{2}(\omega).$ (9)
@@
$\text{Int}=\int_{-\infty}^{\infty}d\omega A^{*}(\omega)^{*}e^{-i\omega t}A(\omega)e^{i\omega(t+\Delta t)}=\int_{-\infty}^{\infty}d\omega|A(\omega)|^{2}e^{i\omega\Delta t}.$ (10)
@@
$P_{res}=|A_{1}|^{2}+|A_{2}|^{2}+2\,\text{Int}\cdot|A_{1}||A_{2}|\cos\phi,$ (12)
where Int is given by expression (11). We see that if $\Delta t\gg t_{\text{coh}}$, the resulting intensity is simply the sum of the intensities of the two sources. This case is referred to as the non-coherent regime.
@@
Now that we have outlined the scheme of coherence damping between separated beams of light (11), let us return to Figs. (1, 2) and consider how different this picture is from heat bath theories, in which decoherence arises from interaction with the environment. In the scheme described here, coherence disappears without any interaction with a medium. Non-coherence results solely from the difference in the travel distances between the two parts of the beam. Nothing happens to the beam as it propagates through a nonabsorbing medium. Moreover, if we let the two separated parts of the light travel an arbitrarily long but equal optical distance before recombining them, they will remain mutually coherent.
Moreover, in a real experiment conducted in Ref. [37], there was an interaction that can be considered as an interaction between the investigated system with the thermal environment. It is an interaction between light and mirrors, and a sufficient portion of light was absorbed by mirrors (gray rectangles in Fig. (1). Due to losses resulting from imperfect reflection at the mirrors, we do not obtain the same intensity at the output as at the input. Therefore, the whole process is not completely reversible. The experiment demonstrated the reversibility of transmission and reflection of coherent light beams at the air–glass interface, and the non-reversibility of the same process for non-coherent light beams. In both cases, the interaction with the mirrors was equally strong, meaning that the same portion of light was absorbed. This interaction, however, did not affect the outcome of the experiment: coherent light beams remained coherent and, under appropriate phase conditions, were successfully recombined into a single beam.
@@
III Non-coherent evolution
@@
As the previous section shows, in the discrete case, the non-coherence condition, $\Delta t\gg t_{\text{coh}}$, converts a reversible unitary transformation into an irreversible stochastic transformation. In the continuous case, unitary evolution is described by the exponential of a skew-Hermitian matrix [41]. In the context of quantum theory, we typically express this as $\mathcal{U}(t)=\exp(i\mathcal{H}t)$, where $\mathcal{H}$ is a Hermitian matrix, particularly representing the Hamiltonian of the system. The continuous analog of a bistochastic transformation is a symmetric continuous-time Markov chain (CTMC). The time dependence of transition probabilities is given by $\mathcal{T}(t)=\exp(\mathcal{Q}t)$, where $\mathcal{Q}$ is the transition rate matrix (Q-matrix) [42]. Its non-diagonal elements are positive numbers representing the rates of transition between states. The diagonal elements $Q_{\alpha\alpha}$ represent the total departure rate from the state $\alpha$. They are negative, such that the sums of elements in each column are zero [^1] In probability texts, the state of the system is often represented by a row, and the transformation matrix is placed on the right of it. We use the convention adopted in physical texts: to represent the state as a column and place the transformation matrix to the left of it. This convention results in a difference in definition: in probability textbooks, the sums of elements in each row of a Q-matrix are zero.. A general Q-matrix serves as the continuous analog of a stochastic transformation, while the analog of a bistochastic transformation is a Q-matrix symmetric over the main diagonal, i.e., $Q_{\alpha\beta}=Q_{\beta\alpha}$. We are going to demonstrate that just as the unitary transformation converts into a bistochastic transformation under the condition of non-coherence, unitary evolution converts into symmetric stochastic evolution.
Figure 3: The diagram illustrates the characteristic scales of the system. The coherence length $l_{\text{coh}}$ is much smaller than the size of the system. For example, we depict the interaction of two bosonic particles resulting in their merging into a single bosonic particle. The regions of the system from which the two particles originate are separated by a distance much greater than the coherence length, and their interaction is therefore non-coherent.
Let us describe a physical system to which our method will be applied. First, we consider a many-particle system described by a $\psi$-function, whose evolution is governed by the Schrödinger equation. We assume the system to be isolated and localized within a finite region of space. For the system to be of interest, its Hamiltonian must include interactions between the constituent particles.
Suppose the system has fixed energy $E$. In that case, the time dependence of its $\psi$-function is given by the exponent $e^{iEt/\hbar}$, and since the amplitudes of the wave function remain constant, the probability distribution is constant as well. However, in our treatment we want to model a non-trivial evolution of probabilities. We assume that, just as the light beam in the previous section did not consist of a single frequency but rather a collection of frequencies within some interval $\omega_{\text{idth}}$, the system under consideration does not have an exact energy $E$, but rather a collection of energies within an interval $\hbar\omega_{\text{idth}}$ around $E$. Since there is a spectral width, there is also a coherence time, defined as $t_{\text{coh}}=1/\omega_{\text{idth}}$ (exactly as in the previous paragraph). Finally, if there is a coherence time, there is also a coherence length, $l_{\text{coh}}=vt_{\text{coh}}$, although these lengths differ for different velocities.
Now we want our system to be sufficiently large, macroscopic, and here is what we mean by this. First, a closed quantum system has a discrete spectrum, and we will make substantial use of this fact in our treatment. However, we assume that the spectral width is much greater than the typical interval between energy levels of the system. Next, we require the size of the system to be much greater than the coherence lengths corresponding to all velocities of interest.
Ideally, we would like to prove that the evolution of such a system already possesses the desired properties: namely, that it is irreversible and that its limiting state corresponds to an equidistribution of probabilities over all microstates of the system. However, the methods we have in mind do not permit a solution in full generality. Therefore, we introduce two additional simplifying assumptions. First, we assume that the interaction between particles is weak. Second, we assume that the system is homogeneous. Both of these assumptions will be clarified further in the text.
”Let us now give a qualitative description of why this system should be irreversible. To this end, we compare it with the light beam investigated in the previous section. There, we subdivided the beam into imaginary segments of length $l_{\text{coh}}$ and found that when photons from different segments interact, the interaction is irreversible (Fig. 2). In the case of the light beam, all photons move in the same direction with the same speed, so in order for photons from different regions to interact, we had to split the beam into two parts with paths of different optical lengths.
In analogy, let us make an imaginative subdivision of the quantum system described earlier into zones of characteristic size $l_{\text{coh}}$. Unlike the beam, the waves that constitute our system propagate in all directions and with different velocities. Consequently, waves originating from different zones naturally cross and interact. However, for our argument to hold, the interaction must be weak enough that a typical interaction involves zones separated by distances much larger than the coherence length (Fig. 3). This is precisely the role of the weak-interaction assumption.
@@
With this qualitative picture in mind, let us now present the mathematical formalism that corresponds to the intuitive idea. We start similarly to a classic simple time-dependent perturbation theory [43] that is aimed to derive Fermi’s golden rule. We seek for a wave function of a many-body system in the form
@@
$\int\limits_{0}^{\Delta t}\exp\left(\frac{i}{\hbar}(E_{k}-E_{l})t\right)A\,dt\approx A\int\limits_{0}^{\Delta t}\exp\left(\frac{i}{\hbar}(E_{k}-E_{l})t\right)dt=-\frac{i\hbar}{E_{k}-E_{l}}\left(e^{\frac{i}{\hbar}(E_{k}-E_{l})\Delta t}-1\right)A.$ (25)
@@
We want to find the appropriate length of this region. We will use $\chi,\overline{\chi}$ in expressions like $\int dE\chi A$, which correspond to the scalar product of amplitude, as a function of energy, over $\chi$. We want norms of $\chi,\overline{\chi}$ be equal: $\int dE|\chi|^{2}=\int dE|\overline{\chi}|^{2}$, which yields
@@
We now want to analyze the non-coherent regime. This means that the state under consideration, $|k\rangle$, belongs to a sheaf of states with a width in the frequency region $\omega_{\text{idth}}$, and we choose a working timescale, $\Delta t$, such that it is much larger than the coherence time. Equivalently, we have $\Delta E\ll\omega_{\text{idth}}$ (Fig. 4b). The quantity $\Delta E$ fits into $\omega_{\text{idth}}$ many times, each time with a different phase difference for each pair of amplitudes, so we can average over the phase. This allows the formula for a series of non-coherent transformations, given in (20), to be applicable. We note that this approximation is in close analogy with the Born-Markov approximation of the Lindblad equation [16], where it is assumed that the heat bath relaxation time is small compared to the relaxation time of the system.
@@
We begin by utilizing an approximation of a continuous spectrum. We can do it since, as we have assumed at the beginning of the section, the spectral width of the system is much greater than the energy difference between the levels of the system. Near a given state $l$, we divide the quantum numbers into two components. The first is $l_{T}$, such that small changes in its value leave the energy unchanged, so it is tangential to the energy surface. The second is $l_{N}$, where a change results in a change in energy. We express energy as a function of the quantum numbers in the form $E(l)=E(l_{0})+l_{N},\mathrm{grad}_{l}E$. Additionally, we recall expression (27) and substitute $\Delta t=\pi\hbar/\Delta E$. We then obtain:
@@
$\frac{1}{\hbar^{2}}\sum_{l}|V_{kl}|^{2}\Delta tP_{l}=\frac{2\pi}{\hbar}C\int dl\frac{\overline{\chi}_{\Delta E}}{\Delta E}|V_{kl}|^{2}P_{l},$ (48)